OpenStudy (anonymous):

ILL GIVE U A MEDAL AND BECOME A FAN!! HELP ME PLEZ!! Use the point-slope form linear equation: y+4=-4(x-2) What is the equation in standard form of a parallel line that passes through (0,-2)?

OpenStudy (anonymous):

are u able to help me? @Cosmichaotic

OpenStudy (anonymous):

Oh yes, sure! Sorry, was afk

OpenStudy (anonymous):

Ok, so just a little quick lesson if you want my help!

OpenStudy (anonymous):

There are 3 forms we can write linear equations in. One is called the Slope Intercept Form. This looks like this: y = mx + b Another one is called Point-Slope Form. This looks like this: y-y1=m(x - x1) This is the form your equation is in. The last one is called Standard Form. This looks like this: Ax + Bx + C = 0

OpenStudy (anonymous):

Fix the last one: Standard Form is: Ax + By + C = 0

OpenStudy (anonymous):

Now. We need to turn your Point-Slope equation into Standard Form.

OpenStudy (anonymous):

Let me know you are with me so far.

OpenStudy (anonymous):

okay. I see where ur going

OpenStudy (anonymous):

Alright cool! So your equation is: \[y+4=-4(x-2)\] Let's get that into Standard Form First. Let's do all the operations we need to to get this thing equal to 0.